OpenStudy (anonymous):

Solve: 1-sin(2πt)=0, for 0≤t≤2.

OpenStudy (michele_laino):

we have: \[\Large \sin \left( {2\pi t} \right) = 1\]

OpenStudy (michele_laino):

now: sin x= 1 when: \[x = \frac{\pi }{2},2\pi + \frac{\pi }{2},4\pi + \frac{\pi }{2}...\]

OpenStudy (michele_laino):

for example, for the first solution, pi/2, we can write: \[2\pi t = \frac{\pi }{2}\] please solve for t

OpenStudy (anonymous):

t=2π?

OpenStudy (triciaal):

no try again

OpenStudy (anonymous):

t=π^2?

OpenStudy (triciaal):